# EraseDiff: Erasing Data Influence in Diffusion Models

### 11 Jan 2024 v1

@@ 3.1 Training objective
1:Well-trained model $\epsilon_{{\bm{\theta}}_{0}}$, forgetting data ${\mathcal{D}}_{f}$ and subset of remaining data ${\mathcal{D}}_{rs}\subset{\mathcal{D}}_{r}$, outer iteration number $S$ and inner iteration number $K$, learning rate $\zeta$ and hyparameter $\lambda$.
2:Parameters ${\bm{\theta}}^{*}$ for the scrubbed model.
3:for iteration $s$ in $S$ do
4:     $\bm{\phi}_{s}^{0}={\bm{\theta}}_{s}$.
5:     Get $\bm{\phi}_{s}^{K}$ by $K$ steps of gradient descent on $f(\bm{\phi}_{s},{\mathcal{D}}_{f})$ start from $\bm{\phi}_{s}^{0}$ using Eq. 8.
6:     Set $\hat{f}(\bm{\phi}_{s},{\mathcal{D}}_{f})=f(\bm{\phi}_{s},{\mathcal{D}}_{f})-f(\bm{\phi}_{s}^{K},{\mathcal{D}}_{f})$.
7:     Update the model: ${\bm{\theta}}_{s+1}={\bm{\theta}}_{s}-\zeta(\nabla_{{\bm{\theta}}_{s}}{\mathcal{F}}({\bm{\theta}}_{s},{\mathcal{D}}_{rs})+\lambda\nabla_{\bm{\phi}_{s}}\hat{f}(\bm{\phi}_{s},{\mathcal{D}}_{f}))$.
8:end for
Eq. 5 could be viewed as a bilevel optimization (BO) problem, where the goal is to minimize an outer objective ${\mathcal{F}}({\bm{\theta}},{\mathcal{D}}_{r})$ whose variables include the solution of another minimization problem w.r.t. an inner objective $f(\bm{\phi},{\mathcal{D}}_{f})$. BO problem is challenging, and most existing methods for BO require expensive manipulation of the Hessian matrix (Liu et al., 2022b). We adopt a practical and efficient algorithm that depends only on first-order gradient information proposed by Liu et al. (2022a) to solve Eq. 5. Our method is shown in Algorithm 1 (Detailed version can be found in Algorithm 2 in Appendix A.1). The objective shown in Eq. 5 will be reformulated as a single-level constrained optimization problem
$\min_{{\bm{\theta}}}{\mathcal{F}}({\bm{\theta}},{\mathcal{D}}_{r}),\hskip 10.00002pt\text{s.t.}\hskip 5.0pt\hat{f}(\bm{\phi},{\mathcal{D}}_{f}):=f(\bm{\phi},{\mathcal{D}}_{f})-f(\bm{\phi}^{*}({\bm{\theta}}),{\mathcal{D}}_{f})\leq 0,$ (7)
where $\bm{\phi}^{*}({\bm{\theta}}):=\text{Alg}({\bm{\theta}},{\mathcal{D}}_{f})=\operatorname*{arg\,min}_{\bm{\phi}}f(\bm{\phi},{\mathcal{D}}_{f})$ and Eq. 7 yields first-order algorithms for non-convex functions. Specifically, $\bm{\phi}^{*}({\bm{\theta}})$ is approximated by running $K$ steps of gradient descent of $f(\bm{\phi},{\mathcal{D}}_{f})$ over ${\mathcal{D}}_{f}$, we set $\bm{\phi}^{0}={\bm{\theta}}$, so
$\bm{\phi}^{k}\equiv\text{Alg}({\bm{\theta}},{\mathcal{D}}_{f})=\bm{\phi}^{k-1}-\zeta_{f}\nabla_{\bm{\phi}^{k-1}}f(\bm{\phi}^{k-1},{\mathcal{D}}_{f}),k=1,\dots,K,$ (8)
where $\zeta_{f}=\zeta$ is the learning rate. We could obtain the approximation $\hat{f}(\bm{\phi},{\mathcal{D}}_{f})=f(\bm{\phi},{\mathcal{D}}_{f})-f(\bm{\phi}^{K},{\mathcal{D}}_{f})$. Then, with $\bm{\phi}={\bm{\theta}}$ and the subset ${\mathcal{D}}_{rs}\subset{\mathcal{D}}_{r}$, the model will be updated via
${\bm{\theta}}^{*}={\bm{\theta}}-\zeta(\nabla_{{\bm{\theta}}}{\mathcal{F}}({\bm{\theta}},{\mathcal{D}}_{rs})+\lambda\nabla_{\bm{\phi}}\hat{f}(\bm{\phi},{\mathcal{D}}_{f})).$ (9)
@@
We use $\lambda=0.1$ by default. $\lambda$ could also be automatically computed for different iterations as shown in the study (Liu et al., 2022a) and results can be found in Appendix A.2.
@@ 4 Related Work
Memorization in generative models. Privacy of generative models has been studied a lot for GANs (Feng et al., 2021; Meehan et al., 2020; Webster et al., 2021) and generative language models (Carlini et al., 2022; 2021; Jagielski et al., 2022; Tirumala et al., 2022; Lee et al., 2023). These generative models often risk replication from their training data. Recently, several studies (Carlini et al., 2023; Somepalli et al., 2023b; a; Vyas et al., 2023) investigated these data replication behaviors in diffusion models, raising concerns about the privacy and copyright issues. Possible mitigation strategies are deduplicating training data and randomizing conditional information (Somepalli et al., 2023b; a), or training models with differential privacy (DP) (Abadi et al., 2016; Dwork et al., 2006; Dwork, 2008; Dockhorn et al., 2022). However, Carlini et al. (2023) show that deduplication is not a perfect solution, and leveraging DP-SGD (Abadi et al., 2016) may cause the training to diverge.
@@
Malicious misuse. Diffusion models usually use training data from varied open sources and when such unfiltered data is employed, there is a risk of it being tainted(Chen et al., 2023b) or manipulated (Rando et al., 2022), resulting in inappropriate generation (Schramowski et al., 2023). They also risk the imitation of copyrighted content, such as mimicking the artistic style (Gandikota et al., 2023a; Shan et al., 2023). To counter inappropriate generation, data censoring (Gandhi et al., 2020; Birhane & Prabhu, 2021; Nichol et al., 2021; Schramowski et al., 2022) where excluding black-listed images before training, and safety guidance where diffusion models will be updated away from the inappropriate/undesired concept (Gandikota et al., 2023a; Schramowski et al., 2023) are proposed. Shan et al. (2023) propose protecting artistic style by adding barely perceptible perturbations into the artworks before public release. Yet, Rando et al. (2022) argue that filtering can still generate disturbing content that bypasses the filter. Chen et al. (2023b) highlight the susceptibility of diffusion models to poison attacks, where target images will be generated with specific triggers.
@@
Machine unlearning. Removing data directly involves retraining the model from scratch, which is inefficient and impractical. Thus, to reduce the computational overhead, efficient machines unlearning methods Romero et al. (2007); Karasuyama & Takeuchi (2010); Cao & Yang (2015); Ginart et al. (2019); Bourtoule et al. (2021); Wu et al. (2020); Guo et al. (2020); Golatkar et al. (2020); Mehta et al. (2022); Sekhari et al. (2021); Chen et al. (2023a); Tarun et al. (2023b) have been proposed. Gandikota et al. (2023a); Heng & Soh (2023); Gandikota et al. (2023b) recently introduce unlearning in diffusion models. Gandikota et al. (2023a; b) mainly focuses on text-to-image models and high-level visual concept erasure. Heng & Soh (2023) adopt Elastic Weight Consolidation (EWC) and Generative Replay (GR) from continual learning to perform unlearning effectively without access to the training data. In this work, we present an unlearning algorithm for diffusion-based generative models. Specifically, we update the model parameters where the outer objective is to maintain utility for the remaining data ${\mathcal{D}}_{r}$, and the inner objective is to erase information about the forgetting data ${\mathcal{D}}_{f}$. Note that both the inner and outer objectives in our formulation and MAML (Rajeswaran et al., 2019) are directed at optimizing the model parameters. Our formulation distinction from MAML lies in its focus: we are not seeking a model adaptable to unlearning but one that effectively erases the influence of data points on the model. Our method allows for easily equipping any plug-and-play loss function(e.g.(Ruiz et al., 2023; Song et al., 2023)) into the unlearning procedure. This is as easy as changing the outer loop to incorporate plug-and-play loss functions if they are deemed to improve the quality of generation.
@@ 5 Experiment
We evaluate the proposed unlearning method EraseDiff in various scenarios, including removing images with specific classes/races/attributes, to answer the following research questions (RQs):
@@
How does the proposed method perform on the public well-trained models from Hugging Face [^2] https://huggingface.co/models?
@@ 5.1 Setup
Experiments are reported on CIFAR10 (Krizhevsky et al., 2009), UTKFace (Zhang et al., 2017), CelebA (Liu et al., 2015), and CelebA-HQ (Lee et al., 2020) datasets. By default, our scrubbed models use $\sim 8K$ images from ${\mathcal{D}}_{r}$ during the unlearning process. Implementation details and more results (e.g., concept/attribute unlearning on CelebA and CelebA-HQ, class unlearning on CIFAR-10 and UTKFace), including the ablation study (ie., the hyper-parameter $\lambda$ that control the balance between ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{f}$, the number of images from ${\mathcal{D}}_{r}$), can be found in the Appendix.
@@
Baselines. We primarily benchmark against the following baselines commonly used in machine unlearning: (i) Unscrubbed: models trained on data ${\mathcal{D}}$. Unlearning algorithms should scrub information from its parameters. (ii) Retrain: models obtained by retraining from scratch on the remaining data ${\mathcal{D}}_{r}$. (iii) Finetune (Golatkar et al., 2020): finetuning models on the remaining data ${\mathcal{D}}_{r}$, ie., catastrophic forgetting. (iv) NegGrad (Golatkar et al., 2020): gradient ascent on the forgetting data ${\mathcal{D}}_{f}$. (v) BlindSpot (Tarun et al., 2023b): the state-of-the-art unlearning algorithm for regression. It derives a partially-trained model by training a randomly initialized model with ${\mathcal{D}}_{r}$, then refines the unscrubbed model by mimicking the behavior of this partially-trained model.
@@
Metrics. Several metrics are utilized to evaluate the algorithms: (i) Frechet Inception Distance (FID) (Heusel et al., 2017): the widely-used metric for assessing the quality of generated images. (ii) Accuracy: the classification rate of a pre-trained classifier, with a ResNet architecture (He et al., 2016) used to classify generated images conditioned on the forgetting classes. A lower classification value indicates superior unlearning performance. (iii) Membership inference attack (MIA): a standard metric for verifying unlearning effect in classification tasks. However, due to the high-quality image generation of diffusion models, which closely resemble real images, MIA might not distinctly differentiate between training and unseen images. (iv) Kullback–Leibler (KL) divergence: distance between the approximator’s output ${\bm{\epsilon}}_{T}$ distribution and the standard Gaussian noise ${\bm{\epsilon}}$ distribution. (v) Weight Distance (WD) (Tarun et al., 2023a): distance between the Retrain models’ weights and other scrubbed models’ weights. WD gives additional insights about the amount of information remaining in the models about the forgetting data.
Table 1: FID score over forgetting classes on conditional DDIM. FID score is evaluated on 50K generated images (5K per class for CIFAR10, 12.5K per class for UTKFace). ‘Accuracy’: the classification accuracy on the corresponding generated images. The pre-trained classifiers have a classification accuracy of 0.81 on CIFAR10 and 0.84 on UTKFace. The generated images conditioned on the forgetting classes significantly deviate from the corresponding real counterparts.
CIFAR10 UTKFace

### 05 Feb 2024 v2

@@ 3.2 Final objective and solution
0:  Well-trained model $\epsilon_{{\bm{\theta}}_{0}}$, forgetting data ${\mathcal{D}}_{f}$ and subset of remaining data ${\mathcal{D}}_{rs}\subset{\mathcal{D}}_{r}$, outer iteration number $E$ and inner iteration number $K$, learning rate $\eta$.
0:  Parameters ${\bm{\theta}}^{*}$ for the scrubbed model.
1:  for iteration $k$ in $E$ do
2:     $\bm{\phi}^{0}={\bm{\theta}}_{k}$.
3:     Get $\bm{\phi}^{K}$ by $K$ steps of gradient descent on $f(\bm{\phi},{\mathcal{D}}_{f})$ starting from $\bm{\phi}^{0}$.
4:     Set $g({\bm{\theta}}_{k})=f({\bm{\theta}}_{k},{\mathcal{D}}_{f})-f(\bm{\phi}^{K},{\mathcal{D}}_{f})$.
5:     Update the model: ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\eta(\nabla_{{\bm{\theta}}_{k}}{\mathcal{L}}({\bm{\theta}}_{k},{\mathcal{D}}_{rs})+\lambda_{k}\nabla_{{\bm{\theta}}_{k}}g({\bm{\theta}}_{k},\bm{\phi}^{K}))$,
6:     where $\lambda_{k}=\max\left\{0,\frac{a_{k}-\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{T}\nabla_{{\bm{\theta}}}{\mathcal{L}}\left(\bm{{\bm{\theta}}_{k}},{\mathcal{D}}_{r}\right)}{\|\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)\|_{2}^{2}}\right\}$.
7:  end for
@@
[…]
Remark.
We set $\lambda_{k}=0.1,\forall k$ by default. $\lambda_{k}$ could also be automatically computed as shown in Algorithm 1 and results can be found in Appendix.
We can characterize the solution of our algorithm as follows:
Theorem 2 (Pareto optimality).
The stationary point obtained by our algorithm is Pareto optimal of the problem $\min_{\bm{\theta}}\left[\mathcal{L}\left(\bm{\theta},\mathcal{D}_{r}\right),f\left(\bm{\theta},\mathcal{D}_{f}\right)\right]$.
Proof.
Let ${\bm{\theta}}^{\ast}$ be the solution to our problem. Recall that for the current ${\bm{\theta}}$, we find $\bm{\phi}^{K}$ to minimize $g({\bm{\theta}},\bm{\phi})=f({\bm{\theta}},{\mathcal{D}}_{f})-f(\bm{\phi},{\mathcal{D}}_{f})$. Assume that we can update in sufficient number of steps $K$ so that $\bm{\phi}^{K}=\bm{\phi}^{*}({\bm{\theta}})=\operatorname{argmin}_{\bm{\phi}\mid{\bm{\theta}}}g({\bm{\theta}},\bm{\phi})=\operatorname{argmax}_{\bm{\phi}\mid{\bm{\theta}}}f(\bm{\phi},{\mathcal{D}}_{f})$. Here $\bm{\phi}|\theta$ means $\bm{\phi}$ is started from ${\bm{\theta}}$ for its update.
The objective aims to minimize $\mathcal{L}({\bm{\theta}},{\mathcal{D}}_{r})+\lambda g({\bm{\theta}},\bm{\phi}^{*}({\bm{\theta}}))$ whose optimal solution is ${\bm{\theta}}^{\ast}$. Note that $g({\bm{\theta}},\bm{\phi}^{*}({\bm{\theta}}))\geq 0$ and it decreases to 00 for minimizing the above sum. Therefore, $g({\bm{\theta}}^{*},\bm{\phi}^{*}({\bm{\theta}}^{*}))=0$. This further means that $f({\bm{\theta}}^{*},{\mathcal{D}}_{f})=f(\bm{\phi}({\bm{\theta}}^{*}),{\mathcal{D}}_{f})$, meaning that ${\bm{\theta}}^{*}$ is the current optimal solution of $f({\bm{\theta}},D_{f})$ because we cannot update further the optimal solution. Moreover, we have ${\bm{\theta}}^{*}$ as the local minima of $\mathcal{L}\left({\bm{\theta}},{\mathcal{D}}_{r}\right)$ because $g\left({\bm{\theta}},\bm{\phi}^{*}({\bm{\theta}}^{*})\right)=0$ and we consider a sufficiently small vicinity around ${\bm{\theta}}^{*}$. ∎
@@ 4 Related Work
Memorization in generative models. Privacy of generative models has been studied a lot for GANs Feng et al. (2021); Meehan et al. (2020); Webster et al. (2021) and generative language models Carlini et al. (2022; 2021); Jagielski et al. (2022); Tirumala et al. (2022); Lee et al. (2023). These generative models often risk replication from their training data. Recently, several studies Carlini et al. (2023); Somepalli et al. (2023b; a); Vyas et al. (2023) investigated these data replication behaviors in diffusion models, raising concerns about the privacy and copyright issues. Possible mitigation strategies are deduplicating training data and randomizing conditional information Somepalli et al. (2023b; a), or training models with differential privacy (DP) Abadi et al. (2016); Dwork et al. (2006); Dwork (2008); Dockhorn et al. (2022). However,  Carlini et al. (2023) shows that deduplication is not a perfect solution, and leveraging DP-SGD Abadi et al. (2016) may cause the training to diverge.
@@
Malicious misuse. Diffusion models usually use training data from varied open sources and when such unfiltered data is employed, there is a risk of it being taintedChen et al. (2023b) or manipulated Rando et al. (2022), resulting in inappropriate generation Schramowski et al. (2023). They also risk the imitation of copyrighted content, e.g., mimicking the artistic style Gandikota et al. (2023a); Shan et al. (2023). To counter inappropriate generation, data censoring Gandhi et al. (2020); Birhane & Prabhu (2021); Nichol et al. (2021); Schramowski et al. (2022) where excluding black-listed images before training, and safety guidance where diffusion models will be updated away from the inappropriate/undesired concept Gandikota et al. (2023a); Schramowski et al. (2023) are proposed.  Shan et al. (2023) propose protecting artistic style by adding barely perceptible perturbations to the artworks before public release. Yet,  Rando et al. (2022) argue that DMs can still generate disturbing content that bypasses the filter.  Chen et al. (2023b) highlight the susceptibility of DMs to poison attacks, where target images are generated with specific triggers.
@@
Machine unlearning. Removing data directly involves retraining the model from scratch, which is inefficient and impractical. Thus, to reduce the computational overhead, efficient machines unlearning methods Romero et al. (2007); Karasuyama & Takeuchi (2010); Cao & Yang (2015); Ginart et al. (2019); Bourtoule et al. (2021); Wu et al. (2020); Guo et al. (2020); Golatkar et al. (2020); Mehta et al. (2022); Sekhari et al. (2021); Chen et al. (2023a); Tarun et al. (2023b) have been proposed. Several studies Gandikota et al. (2023a; b); Heng & Soh (2023a; b); Fan et al. (2023); Zhang et al. (2023) recently introduce unlearning in diffusion models. Most of them Gandikota et al. (2023a; b); Heng & Soh (2023a); Zhang et al. (2023) mainly focus on text-to-image models and high-level visual concept erasure.  Heng & Soh (2023b) adopt Elastic Weight Consolidation (EWC) and Generative Replay (GR) from continual learning to perform unlearning effectively without access to the training data. Heng and Soh’s method can be applied to a wide range of generative models, however, it needs the computation of FIM for different datasets and models, which may lead to significant computational demands.
@@ 5 Experiment
We evaluate EraseDiff in various scenarios, including removing images with specific classes/concepts, to answer the following research questions (RQs):
@@
How does the proposed method perform on the public well-trained models?
@@ 5.1 Setup
Experiments are reported on CIFAR10 Krizhevsky et al. (2009) with DDPM and DDIM, Imagenette Howard & Gugger (2020) with Stable Diffusion (SD) for class-wise forgetting, CelebA-HQ Lee et al. (2020) with DDPM and I2P Schramowski et al. (2023) dataset with SD for concept-wise forgetting. For all SD experiments, we use the open-source SD v1.4 Rombach et al. (2022) checkpoint as the pre-trained model. We fine-tune the cross-attention (denoted as EraseDiff-x) and the unconditional weights (denoted as EraseDiff-u) of the U-Net module in SD. In the experiment on forgetting the ‘airplane’ class with DDPM and nudity erasure with SD, we have no access to the training data. Implementation details and more results (e.g., the ablation study (e.g., the hyper-parameter $\lambda$ that controls the balance between ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{f}$), can be found in the Appendix.
@@
Baselines. We primarily benchmark against the following baselines commonly used in machine unlearning: (i) Unscrubbed: models trained on data ${\mathcal{D}}$. Unlearning algorithms should scrub information from its parameters. (ii) Retrain: models obtained by retraining from scratch on the remaining data ${\mathcal{D}}_{r}$. (iii) Finetune Golatkar et al. (2020): finetuning models on the remaining data ${\mathcal{D}}_{r}$, ie., catastrophic forgetting. (iv) NegGrad Golatkar et al. (2020): gradient ascent on the forgetting data ${\mathcal{D}}_{f}$. (v) BlindSpot Tarun et al. (2023b): the state-of-the-art unlearning algorithm for regression. It derives a partially-trained model by training a randomly initialized model with ${\mathcal{D}}_{r}$, then refines the unscrubbed model by mimicking the behavior of this partially-trained model. (vi) ESD Gandikota et al. (2023a): fine-tune the model’s conditional prediction away from the erased concept. (vii) For class-wise forgetting, we compare against Selective Amnesia (SA) Heng & Soh (2023b) when ${\mathcal{D}}_{r}^{{}^{\prime}}$ is generated by Generative Replay (GR). (viii) For nudity-concept unlearning, we also compare against SD v2.1 Rombach et al. (2022) which is trained on a dataset filtered for nudity.
@@
Metrics. Several metrics are utilized to evaluate the algorithms: (i) Frechet Inception Distance (FID) Heusel et al. (2017): the widely-used metric for assessing the quality of generated images. (ii) $P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$ Heng & Soh (2023b): the classification rate of a pre-trained classifier $P_{\psi}({\mathbf{y}}|{\mathbf{x}})$, with a ResNet architecture He et al. (2016) used to classify generated images conditioned on the forgetting classes. A lower classification value indicates superior unlearning performance. (iii) $H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$ Heng & Soh (2023b): the average entropy of the classifier’s output distribution given ${\mathbf{x}}_{f}$, which is $H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))=-\mathbb{E}[\sum_{i}P_{\psi}({\mathbf{y}}=c_{i}|{\mathbf{x}})logP_{\psi}({\mathbf{y}}=c_{i}|{\mathbf{x}})]$. Suppose that all information about $c_{f}$ is erased, the classifier becomes maximally uncertain and the entropy therefore would approach $-\sum_{i=1}^{10}\frac{1}{10}log\frac{1}{10}=2.30$ for CIFAR10. (iv) Weight Distance (WD) Tarun et al. (2023a): distance between the Retrain models’ weights and other scrubbed models’ weights. WD gives additional insights about the amount of information remaining in the models about the forgetting data.
Table 1: Results on CIFAR10 with conditional DDPM. $*{}^{*}$ and EraseDiff (${\mathcal{D}}_{r}^{{}^{\prime}}$) have no access to the training data. $H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$ and $P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$ indicate the entropy of the classifier’s distribution and the probability of the forgotten class, respectively Heng & Soh (2023b). Note that $*{}^{*}$ is trained for 20K training steps, while ours is trained only for 200 training steps.
Unscrubbed EraseDiff EraseDiff (${\mathcal{D}}_{r}^{{}^{\prime}}$) $*{}^{*}$
FID $\downarrow$ 9.69 7.97 11.74 9.08
Precision $\uparrow$ 0.392 0.43 0.379 0.412
Recall $\uparrow$ 0.787 0.740 0.76 0.767
$H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$ 0.037 2.017 1.59 1.47
$P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$ 0.972 0.216 0.355 0.156
Table 2: Computational overhead. Time (s) is the average duration of one training step, measured over five runs.
Method Compute Memory (MiB) Time (s)
SA ${\mathcal{O}}(N_{r}n+Nl(n))$ 3352.3 9.52
EraseDiff ${\mathcal{O}}(Nl(n))$ 3360.3 4.16
Figure 2: Generated images (forgetting the ‘airplane’ class).
Table 3: Results on CIFAR10 with conditional DDIM. Although NegGrad and BlindSpot could scrub the information of the forgetting classes, they cannot maintain the model utility. Other methods fail to scrub the relevant information completely.